sin x+sin y = ¼ and sin x- sin y= 1/5 then the value of 4.cot(x-y/2) a.5cot(x-y/2) b.5tan(x-y/2) c.5cot(x+y/2) d.5tan(x+y/2) this question is correct please find the value..please

sin x+sin y =  ¼  and sin x- sin y= 1/5  then the value of 4.cot(x-y/2)
a.5cot(x-y/2)
b.5tan(x-y/2)
c.5cot(x+y/2)
d.5tan(x+y/2)
this question is correct please find the value..please

Grade:11

3 Answers

Arun
5 years ago
Dear student
 
2sin x + y/2  cos x-y/2  = ¼
 
and 
2 cos x+y/2  sinx-y/2   = 1/5
 
now divide
 
tan x + y/2 cotx-y/2  = 5/4
 
4 cot x-y/2 = 5 cot x+y/2
 
option C is correct
Susmita
5 years ago
sinx+siny=1/4
Or,2sin(x+y/2)cos(x-y/2)=1/4........(eq1)
sinx-siny=1/5
Or,2cos(x+y/2)sin(x-y/2)=1/5........(eq2)
Divide eq1 by eq2.
tan(x+y/2)cot(x-y/2)=5/4
Or,4cot(x-y/2)=5cot(x+y/2)
 
SANDEEP
2 years ago
Dear student
2sin x + y/2  cos x-y/2  = ¼
2 cos x+y/2  sinx-y/2   = 1/5
now divide
tan x + y/2 cotx-y/2  = 5/4
 
4 cot x-y/2 = 5 cot x+y/2
 
option C is correct